# appendixtocaptAppendix

### v3

@@ 2.4 Matrix Squaring, Learning, and Continuation
Let $P_{1}$ be the one-step transition matrix for the random walk on the discrete lattice. $P_{1}$ depends on obstacles in the environment and amounts to local one-step exploration, as detailed in subsection 2.1.
@@
We calculate $t$-step transition matrix $P_{t}$ for a discrete set of $t$, $\mathcal{T}=\{t=2^{k},k=1,...,K\}$, via $P_{2t}=P_{t}^{2}$. The matrix squaring is very efficient for calculating $P_{t}$ for $t\in\mathcal{T}$, and these $P_{t}$ correspond to explorations of different spatial scales $\sqrt{t}$, where the adjacent spatial scales in $\mathcal{T}$ have a ratio $\sqrt{2}$, mirroring grid cell module scaling ($\sim$1.4–1.7) [57]. In addition to matrix squaring, one can design any discrete sequence $\mathcal{T}=\{t_{k},k=1,\ldots,K\}$, and calculate $P_{t}$ for $t\in\mathcal{T}$ using matrix multiplication.
@@
We learn a separate population of place cells $h(x,t)$ for each $t\in\mathcal{T}$, by minimizing the least squares error:
$\mathcal{L}=\sum_{x,y}\left[q(y|x,t)-\langle h(x,t),h(y,t)\rangle\right]^{2}$ (9)
@@
We learn $h(x,t)$ over the discrete lattice, where $\sum_{x,y}$ in (9) is over all pairs of points on the lattice. We optimize this objective using the AdamW optimizer [36], where after each iteration, for each $x$, we set the negative elements of $h(x,t)$ to 0, and then normalize $h(x,t)$ so that $\|h(x,t)\|=1$.
In high dimension, random vectors tend to have near zero inner product, so they provide good fit for non-adjacent $x$ and $y$. Learning is mainly to fit adjacent $x$ and $y$.
@@
The learning reduces pairwise adjacency relationships $(q(y|x,t),\forall x,y)$ into individual embeddings $(h(x,t),\forall x)$, which collectively form a map of the environment.
@@
After learning, we can use bi-linear interpolation to make $h(x,t)$ a continuous map over $x$. As a result, $q(y|x,t)=\langle h(x,t),h(y,t)\rangle$ also becomes continuous, approximating the transition kernel of the continuous heat equation. The continuous $q(y|x,t)$ can then elegantly guide planning of smooth trajectories in continuous space instead of discrete lattice.
The non-negative constraint enables us to learn localized bump-like place fields. Without non-negative constraint, the place fields will be wave-like instead of bump-like. The localized place fields make the learned embeddings interpretable as place cells, enabling us the shift attention from the population of cells to individual cells. The localized place fields collectively form a tiling of the environment.
@@
2.5 Straight-Forward Path Planning via Adaptive Gradient Following
@@
We use adaptive gradient following for goal-directed navigation. When navigating from a current position $x$ to a target location $y$, we select the next position $x_{\text{next}}$ from the neighborhood $\partial(x)$ where $\partial(x)=\{z:z=x+\Delta r(\cos\theta,\sin\theta)\}$, $\Delta r$ is the step size, and $\theta$ is discretized into equally spaced direction in $[0,2\pi)$. We use $\partial(x)$ for continuous $x$ for path planning to differentiate from $N(x)$ in the discrete lattice for random walk.
@@
  • Compute the gradient of the normalized transition probability for each neighbor $z\in\partial(x)$ and each scale $t$:
    $\Delta(z,t)$ $=q(y|z,t)-q(y|x,t)$ (10)
    $=\langle h(y,t),h(z,t)\rangle-\langle h(y,t),h(x,t)\rangle$ (11)
    $=\frac{1}{2}\left[\|h(x,t)-h(y,t)\|^{2}-\|h(z,t)-h(y,t)\|^{2}\right]$ (12)
  • Select the scale $t^{*}$ that provides the strongest directional signal:
    $t^{*}=\operatorname*{arg\,max}_{t\in\mathcal{T}}\max_{z\in\partial(x)}\Delta(z,t)$ (13)
@@
    $x_{\text{next}}=\operatorname*{arg\,max}_{z\in\partial(x)}\Delta(z,t^{*})$ (14)
Note that in the above algorithm, $x$ and $z$ are continuous, because $h(x,t)$ is made continuous in $x$ with bi-linear interpolation after learning on discrete lattice. This allows planning smooth paths in continuous domain.
$\Delta(z,t)$ measures the reduction in the angle or squared Euclidean distance. The path planning seeks maximal reduction in angle or Euclidean distance. Therefore we call it the straight-forward path planning in the embedding space, where the vector $h(x,t)$ rotates straightly to $h(y,t)$ on the path.
This approach selects the scale $t^{*}$ that provides the clearest guidance for the current navigation step. Intuitively, larger scales provide better guidance for distant goals, while smaller scales offer more precise navigation for nearby goals. The adaptive scale selection mechanism automatically finds this optimal scale at each step, similar to choosing the most appropriate “ruler” for measuring at the current distance.
Now consider the idealized continuous limit where $\Delta r\rightarrow 0$, $\theta$ is continuous, and $t$ is continuous. The path planning algorithm follows the gradient
$\nabla_{x}q(y|x,t)=\nabla_{x}q(x|y,t)=\frac{1}{2}\nabla_{x}\|h(x,t)-h(y,t)\|^{2}$ (15)
where $q(y|x,t)=q(x|y,t)$ due to symmetry. The planed trajectory is the gradient ascent flow: $dx(\tau)/d\tau=\nabla_{x}q(x(\tau)|y,t(\tau))$, where $\tau$ is the time on the planed trajectory, $t(\tau)=\arg\max_{t}\|\nabla_{x}q(x(\tau)|y,t)\|$ is the optimal scale at time $\tau$ with the maximal gradient.
For the goal $y$, $q(y|x,t)$, as a function of $x$, has a unique maximum at $x=y$ with $q(y|x=y,t)=1$ for all $t$. The gradient-based navigation framework has the following key properties that ensure computational efficiency and biological plausibility:
@@
    Adaptive scale selection dynamically adjusts $\sqrt{t^{*}}\propto d(x,y)$ for precise navigation in the open environment, mirroring hippocampal spatial tuning [38].
@@
    The smooth gradient field $\nabla_{x}q(x|y,t)$, with a unique maximum of $q(x|y,t)$ at the goal $x=y$, ensures trap-free paths, aligning with hippocampal navigation [38, 41].
@@
    Planned paths match the shortest path for small $t$, but prioritize topological connectivity for large $t$, reflecting cognitive map robustness [41, 59].
@@
    Near obstacles, $\nabla_{x}q(x|y,t)$ flows parallel to boundaries, preventing collisions, akin to hippocampal obstacle avoidance [38].
@@
These properties, detailed in Appendix D, underpin the framework’s alignment with neural mechanisms.
@@ 3 Experiments
We designed experiments to evaluate both the biological plausibility of our position embedding framework and its functional capabilities in path planning. All experiments are conducted in a simulated environment. We first examine the positional representation and population density profiles in an open field. Then we create more complex environments by adding obstacles. For implementation details, see Appendix G.
Figure 1: Place Cell Representations and Navigation in Open Field Environment. (A) Goal-directed path planning trajectories with adaptive temporal scale selection. (B) Normalized transition probability kernels $q(y|x,t)$ at multiple temporal scales with gradient vector fields. (C) Spatial activation profiles of place cells. (D) Collective environmental coverage of overlapping place fields (different colors represent different cells’ activation fields).
@@ 3.1 Place Cell Representations in Open Field
We begin our numerical experiments in a simple open field environment to demonstrate fundamental properties. The environment consists of a $40\times 40$ lattice grid. For the transition kernel, we employ a $3\times 3$ neighborhood, meaning each position has at most 8 neighboring points, with transition probability $p_{\rm move}=1/9$ distributed uniformly among neighbors as well as self. Within this environment, we learn $h(x,t)$ embeddings across all lattice points. For hyperparameters, the total number of place cells is set to $500$. The temporal dynamics are captured across multiple scales using the time parameter $t=2^{k}$, where $k=1,2,...,11$, allowing us to analyze system behavior across spatial scales.
3.1.1 Multi-Time Scale Transitions
To evaluate the normalized transition probability kernel $q(y|x,t)$, we visualize the learned Gaussian-like patterns across multiple temporal scales in Figure 1 (B). Each panel depicts the transition probability distribution for time scales $t=16,64,256$ with position $y$ fixed at the environment center.
The gradient fields reveal a critical temporal dependency: as $t$ increases, gradient magnitudes intensify for locations distant from the center, while maintaining directional integrity toward the target. This multi-scale property is particularly significant for path planning applications, as it enables a multi-resolution navigational framework.
@@ 3.1.2 Spatial Activation Patterns
Our model is trained by minimizing Equation 9. The optimization results demonstrate remarkable fidelity in approximating normalized transition probabilities through inner products of population embedding vectors with correlation coefficients above $0.9$ for all scales. Figure 1 (C) illustrates the learned activation patterns of $h(x,t)$ at two representative scales $t=4$ and $t=64$. The spatial receptive fields of individual units emerge naturally from our training process, exhibiting the characteristic Gaussian-like firing fields centered at specific locations within the open environment. This spatial tuning closely resembles the well-documented properties of hippocampal place cells observed in rodent navigation studies [45, 46].
In Figure 1 (D), we visualize the collective spatial coverage of our learned place cell population by overlaying multiple receptive fields on a single map. We use multiple colors to distinguish different place cell activation fields. For the scale $t=4$, we randomly sample 200 place cells, while for the broader scale $t=64$, we display $20$ cells. The population-level position embeddings show comprehensive spatial tiling across the entire environment, with overlapping receptive fields ensuring complete coverage without significant gaps or under-represented regions. Furthermore, the distribution of place field centers appears relatively uniform across the environment, indicating that our method successfully avoids representational biases toward particular regions.
@@ 3.1.3 Path Planning and Adaptive Scale Selection
To evaluate the navigational capabilities of our model, we implemented the gradient-based path planning. We randomly selected start and target locations within the open field environment. At each step, the agent evaluates potential next positions by sampling orientation $\theta$ from $36$ equally spaced directions within $[0,2\pi)$ and a fixed radius $\Delta r$ of one grid unit. Even though our model is learned in a $40\times 40$ discrete lattice, in path planning, we can reach continuous positions $x$, where the representation of its location $h(x,t)$ is calculated by linear interpolation of the 4 closest neighbors. Movement direction is determined by maximizing the gradient $q(y|z,t)-q(y|x,t)$ where $y$ is the target location, $x$ is the current position, and $z$ represents each candidate’s next position. And for each step, we choose the $z$ with the largest gradient. Our results in Figure 1 (A) demonstrate that the learned model consistently generates near-optimal trajectories.
A key feature of our position embedding model is the adaptive scale selection mechanism, where $t^{*}$ is selected with maximal gradient. The mechanism is illustrated in Figure 1 (A), where different colors represent the selected time scale $t^{*}$ at each navigational step. This reveals a systematic progression from coarse to fine scales as the agent approaches its goal. This pattern emerges naturally from the gradient fields associated with different scales.
When navigating to distant goals, larger scales ($t\geq 128$) are selected, which provides clearer guidance for long-range planning. As the agent reaches medium distances from the goal, the selected scales transition to intermediate values ($t=16$ to $t=64$), balancing directional guidance with increasing spatial resolution. In the final approach phase, the system converges on the smallest available scales ($t=2$ to $t=8$), which offered the most precise local guidance.

### v4

@@ 2.4 Matrix Squaring, Learning, and Continuation
Let $P_{1}$ be the one-step transition matrix for the random walk on a discrete lattice. $P_{1}$ depends on obstacles in the environment and amounts to local one-step exploration, as detailed in subsection 2.1.
@@
We calculate $P_{\tau}$ for a discrete set of $\tau$, $\mathcal{T}=\{\tau=2^{k},k=1,...,K\}$, via $P_{2\tau}=P_{\tau}^{2}$. The matrix squaring is very efficient for calculating $P_{\tau}$ for $\tau\in\mathcal{T}$, and these $P_{\tau}$ correspond to explorations of different spatial scales $\sqrt{\tau}$, where the adjacent spatial scales in $\mathcal{T}$ have a ratio $\sqrt{2}$, mirroring grid cell module scaling ($\sim$1.4–1.7) [22]. In addition to matrix squaring, one can design any discrete sequence $\mathcal{T}=\{t_{k},k=1,\ldots,K\}$, and calculate $P_{\tau}$ for $\tau\in\mathcal{T}$ using matrix multiplication.
@@
We learn a separate population of place cells $h(x,\tau)$ for each $\tau\in\mathcal{T}$, by minimizing the least squares error:
$\mathcal{L}=\sum_{x,y}\left[q(y|x,\tau)-\langle h(x,\tau),h(y,\tau)\rangle\right]^{2}$ (8)
@@
We learn $h(x,\tau)$ over the discrete lattice, where $\sum_{x,y}$ in (8) is over all pairs of points on the lattice. We optimize this objective using the AdamW optimizer [23], where after each iteration, for each $x$, we set the negative elements of $h(x,\tau)$ to 0, and then normalize $h(x,\tau)$ so that $\|h(x,\tau)\|=1$.
@@
The learning reduces pairwise adjacency relationships $(q(y|x,\tau),\forall x,y)$ into individual embeddings $(h(x,\tau),\forall x)$, which collectively form a map of the environment.
@@
After learning, we can use bi-linear interpolation to make $h(x,\tau)$ a continuous map over $x$. As a result, $q(y|x,\tau)=\langle h(x,\tau),h(y,\tau)\rangle$ also becomes continuous, approximating the transition kernel of the continuous heat equation. The continuous $q(y|x,\tau)$ can then elegantly guide path planning in continuous space instead of discrete lattice.
2.5 Euclideanized Cognitive Map
Geometrically, $\langle h(x,\tau),h(y,\tau)\rangle$ is the cosine of the angle between the normalized vectors $h(x,\tau)$ and $h(y,\tau)$. Moreover, we have
$\frac{1}{2}\|h(x,\tau)-h(y,\tau)\|^{2}=1-\langle h(x,\tau),h(y,\tau)\rangle=1-q(y|x,\tau)$ (9)
That is, the angle or the Euclidean distance between $h(x,\tau)$ and $h(y,\tau)$ encodes proximity or adjacency between $x$ and $y$. Thus, our method geometrizes transition probability.
$(h(x,\tau),\forall x)$ is a 2D manifold in the high-dimensional embedding space. This 2D manifold is an embedding of the 2D physical space. Because of the flexibility endowed by the high-dimensional embedding space, the path between $x$ and $y$ on the manifold is essentially “Euclideanized” even though the physical path is far from being Euclidean.
2.6 Emergent Sparsity from Non-negativity and Orthogonality
The conjunction of non-negativity and orthogonality induces emergent sparsity in the representational code. Let $h(x,\tau)\in\mathbb{R}_{+}^{n}$ denote the representation at position $x$ and scale $\sqrt{\tau}$, such that $\langle h(x,\tau),h(y,\tau)\rangle=q(y|x,\tau)$, where $q(y|x,\tau)$ measures spatial adjacency at scale $\sqrt{\tau}$. If two positions $x$ and $y$ are non-adjacent ($q(y|x,\tau)=0$), the corresponding vectors must be orthogonal. Because all components of $h(x,\tau)$ are non-negative, orthogonality implies disjoint support—no component can be active in both representations. Consequently, the joint requirement of non-negativity and orthogonality forces most components to be zero, yielding sparse, localized activity patterns. As a result, each place cell fires at a localized place, and the population of place cells collectively tile the environment. The symbolic representation thus emerges automatically from population-based representation.
The scale parameter $\tau$ modulates this sparsity: for small $\tau$, most position pairs are orthogonal, producing highly localized “place fields”; for large $\tau$, overlaps increase and place fields broaden. Thus, the locality of place-cell activation follows directly from the geometry of non-negative, inner-product spectral embeddings. Appendix F provides technical details.
@@ 2.6 Emergent Sparsity from Non-negativity and Orthogonality
2.7 Straight Forward Path Planning with Adaptive Scale Selection
@@
We use adaptive gradient following for goal-directed navigation. When navigating from a current position $x$ to a target location $y$, we select the next position $x_{\text{next}}$ from the neighborhood $\partial(x)$ where $\partial(x)=\{z:z=x+\Delta r(\cos\theta,\sin\theta)\}$, $\Delta r$ is the step size, and $\theta$ is discretized into equally spaced direction in $0,2\pi)$. We use $\partial(x)$ for continuous $x$ for path planning to differentiate from $N(x)$ in the discrete lattice for random walk.
@@
  • Compute the gradient of the normalized transition probability for each neighbor $z\in\partial(x)$ and each scale $\tau$:
    $\Delta(z,\tau)=q(y|z,\tau)-q(y|x,\tau)=\langle h(y,\tau),h(z,\tau)\rangle-\langle h(y,\tau),h(x,\tau)\rangle$ (10)
  • Select the scale $\tau^{*}$ that provides the strongest directional signal:
    $\tau^{*}=\operatorname*{arg\,max}_{\tau\in\mathcal{T}}\max_{z\in\partial(x)}\Delta(z,\tau)$ (11)
@@
    $x_{\text{next}}=\operatorname*{arg\,max}_{z\in\partial(x)}\Delta(z,\tau^{*})$ (12)
Note that in the above algorithm, $x$ and $z$ are continuous, because $h(x,\tau)$ is made continuous in $x$ with bi-linear interpolation after learning on discrete lattice.
$\Delta(z,\tau)$ measures the reduction in the angle or squared Euclidean distance. The path planning seeks maximal reduction in angle or Euclidean distance. Therefore we call it the straight forward path planning in the Euclideanized embedding space, where the vector $h(x,\tau)$ rotates straightly to $h(y,\tau)$ on the path.
This approach selects the scale $\tau^{*}$ that provides the clearest guidance for the current navigation step. Intuitively, larger scales provide better guidance for distant goals, while smaller scales offer more precise navigation for nearby goals. The adaptive scale selection mechanism automatically finds this optimal scale at each step, similar to choosing the most appropriate “ruler” for measuring at the current distance.
Now consider the idealized continuous limit where $\Delta r\rightarrow 0$, $\theta$ is continuous, and $\tau$ is continuous. The path planning algorithm follows the gradient $\nabla_{x}q(y|x,\tau)=\nabla_{x}q(x|y,\tau)$ where $q(y|x,\tau)=q(x|y,\tau)$ due to symmetry. The planed trajectory is the gradient ascent flow: $dx(t)/dt=\nabla_{x}q(x(t)|y,\tau(t))$, where $t$ is the time on the planed trajectory, $\tau(t)=\arg\max_{\tau}\|\nabla_{x}q(x(t)|y,\tau)\|$ is the optimal scale at time $t$ with the maximal gradient.
The gradient-based navigation framework has the following key properties that ensure computational efficiency and biological plausibility:
@@
    Adaptive scale selection dynamically adjusts $\sqrt{\tau^{*}}\propto d(x,y)$ for precise navigation in the open environment, mirroring hippocampal spatial tuning [11].
@@
    The unimodal smooth gradient field $\nabla_{x}q(x|y,\tau)$, with a unique maximum at the goal, ensures trap-free paths, aligning with hippocampal navigation [11, 9].
@@
    Planned paths match the shortest path for small $\tau$, but prioritize topological connectivity for large $\tau$, reflecting cognitive map robustness [9, 3].
@@
    Near obstacles, $\nabla_{x}q(x|y,\tau)$ flows parallel to boundaries, preventing collisions, akin to hippocampal obstacle avoidance [11].
@@
These properties, detailed in Appendix G, underpin the framework’s alignment with neural mechanisms.
2.8 Theta-Phase Procession Based on Angle-Phase Duality
The hippocampus not only encodes spatial relationships but also organizes temporal dynamics through theta phase precession, where place cells fire at specific phases of the theta rhythm as an animal traverses their fields [24, 25]. We extend our population embedding framework to model this phenomenon.
As explained above, the non-negativity and inner product structure of the position embeddings induces localized sparse patterns, i.e., each place cell only fires at a localized field around a place, and together the place fields of all the place cells tile the environment. Let $\mu_{i}=\arg\max_{x}h_{i}(x,\tau)$, i.e., the center of the place field for cell $i$. The theta phase of cell $i$ can be defined in terms of the angle between $h(x,\tau)$ and $h(\mu_{i},\tau)$. As $x$ approaches $\mu_{i}$, the angle changes from $\pi/2$ to 0, and as $x$ moves away from $\mu_{i}$, the angle changes from 0 to $\pi/2$. Appendix H provides details.
2.9 Integrating Grid Cells
For scientific reductionism, we focus on place cells without incorporating grid cells. A natural extension would explore the relationship $h(x,\tau)=W(\tau)g(x)$, where $g(x)$ represents the vector of grid cell activations and $W(\tau)$ is a learned transformation matrix, potentially unifying our framework with the complementary roles of place and grid cells in spatial representation [9]. Appendix I provides the formulation.
@@ 3 Experiments
We design experiments to evaluate both the biological plausibility of our position embedding framework and its functional capabilities in path planning. All experiments are conducted in a simulated environment. We first examine the positional representation and population density profiles in an open field. Then we create more complex environments by adding obstacles. For implementation details, see Appendix J; For additional experiment visualizations, see Appendix K.
Figure 1: Place Cell Representations and Navigation in Open Field Environment. (A) Goal-directed path planning trajectories with adaptive scale selection (selected scale is color coded, red for big scale and blue for small scale). (B) Normalized transition probability kernels $q(y|x,\tau)$ at multiple scales with gradient vector fields. $y$ is fixed at the center of the environment. (C) Learned activation patterns of $h(x,\tau)$ at different scales across randomly chosen cells, exhibiting Gaussian-like firing fields centered at specific locations within the open environment.
@@ 3.1 Place Cell Representations in Open Field
We begin our numerical experiments in a simple open field environment to demonstrate fundamental properties. The environment consists of a $40\times 40$ lattice grid. For the transition kernel, we employ a $3\times 3$ neighborhood, meaning each position has at most 8 neighboring points, with transition probability $p_{\rm move}=1/9$ distributed uniformly among neighbors as well as self. Within this environment, we learn $h(x,\tau)$ embeddings across all lattice points. For hyperparameters, the total number of place cells is set to $500$. The temporal dynamics are captured across multiple scales using the time parameter $\tau=2^{k}$, where k=1,2,,11k=1,2,...,11, allowing us to analyze system behavior across spatial scales.
3.1.1 Multi-Scale Transitions
To evaluate the normalized transition probability kernel $q(y|x,\tau)$, we visualize the learned Gaussian-like patterns across multiple temporal scales in Figure 1 (B). Each panel depicts the transition probability distribution for time scales $\tau=16,64,256$ with position $y$ fixed at the environment center.
The gradient fields reveal a critical temporal dependency: as $\tau$ increases, gradient magnitudes intensify for locations distant from the center, while maintaining directional integrity toward the target. This multi-scale property is particularly significant for path planning applications, as it enables a multi-resolution navigational framework.
@@ 3.1.2 Spatial Activation Patterns
Our model is trained by minimizing Equation 8. The optimization results demonstrate remarkable fidelity in approximating normalized transition probabilities through inner products of population embedding vectors with correlation coefficients above $0.9$ for all scales. Figure 1 (C) illustrates the learned activation patterns of $h(x,\tau)$ at scales $\tau=4,16,64,256$ for randomly chosen cells. The spatial receptive fields of individual units emerge naturally from our training process, exhibiting the characteristic Gaussian-like firing fields centered at specific locations within the open environment. This spatial tuning closely resembles the well-documented properties of hippocampal place cells observed in rodent navigation studies [1, 2].
@@ 3.1.3 Path Planning and Adaptive Scale Selection
To evaluate the navigational capabilities of our model, we implemented the gradient-based path planning. We randomly selected start and target locations within the open field environment. At each step, the agent evaluates potential next positions by sampling orientation $\theta$ from $36$ equally spaced directions within $0,2\pi)$ and a fixed radius $\Delta r$ of one grid unit. Even though our model is learned in a $40\times 40$ discrete lattice, in path planning, we can reach continuous positions $x$, where the representation of its location $h(x,\tau)$ is calculated by linear interpolation of the 4 closest neighbors. Movement direction is determined by maximizing the gradient $q(y|z,\tau)-q(y|x,\tau)$ where $y$ is the target location, $x$ is the current position, and $z$ represents each candidate next position. And for each step, we choose the $z$ with the largest gradient. Our results in Figure 1 (A) demonstrate that the learned model consistently generates near-optimal trajectories.
A key feature of our position embedding model is the adaptive scale selection mechanism, where $\tau^{*}$ is selected with maximal gradient. The mechanism is illustrated in Figure 1 (A), where different colors represent the selected time scale $\tau^{*}$ at each navigational step. This reveals a systematic progression from coarse to fine scales as the agent approaches its goal. This pattern emerges naturally from the gradient fields associated with different scales.
When navigating to distant goals, larger scales ($\tau\geq 128$) are selected, which provides clearer guidance for long-range planning. As the agent reaches medium distances from the goal, the selected scales transition to intermediate values ($\tau=16$ to $\tau=64$), balancing directional guidance with increasing spatial resolution. In the final approach phase, the system converges on the smallest available scales ($\tau=2$ to $\tau=8$), which offered the most precise local guidance.